# Safe MPC Alignment with Human Directional Feedback

### v2

@@ VI-D Ablation Study
In this session, we conduct two ablation studies to show the performance of our algorithm given different design parameters/aspects. The task is still based on pendulum environment with settings identical to Section VI-A).
@@ VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
In the first study, we compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$ in our algorithm: MVE centering (used in our method) and random sampling. For the latter, $\boldsymbol{\theta}_{i}$ is chosen by sampling from the uniform distribution supported by $\Theta_{i-1}$. We compare the total number of corrections needed at learning convergence under the two strategies.
With 50 independent runs, we observe that the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling requires $11.88\pm 3.50$ corrections. While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance. This is understandable given the sampling randomness. We also observed that sampling process introduced large computational overhead, which is around 25% time slower than MVE centering. In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency in finding $\boldsymbol{\theta}_{i}$. Because selecting $\boldsymbol{\theta}_{i}$ is casted as a convex programming for MVE centering, such advantage would be more significant for high-dimensional parameter space.
@@ VI-D2 Choosing $\gamma$
In the second study, we investigate how the setting of $\gamma$ affects the learning performance. We vary $\gamma$ from 0.01 to 10. Fig. 14 shows the total number of corrections needed at convergence versus different $\gamma$s.
The result in Fig. 14 shows that the learning performance remains the same given different settings of $\gamma$. Since $\gamma$ is weights parameter to control the safety conservativeness in the penalty-based MPC policy in (3), it means our method can successfully learn a safety constraint when the robot executes with safe MPC policy of different conservativeness levels.
Figure 14: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale. Safe MPC alignment can converge with a similar number of corrections (iterations) under different levels of $\gamma$.
@@ VII User Study in Simulation Games
In this section, we test the Safe MPC Alignment method in two real-human user studies using two simulation games developed based on MuJoCo [62]. One is the drone navigation game and the other the Franka robot arm reaching game. In both games, the robot is controlled by an MPC controller and human participants provide real-time corrections through a keyboard. The objective of each game is to teach a robot to successfully learn a safe MPC to move through environments with unknown obstacles. In following, Subsection VII-A describes the basic setting of the two games. Subsection VII-B details the user study procedure and the recruitment of human participants. The results and analysis are given in VII-C.
VII-A Setting of Two Simulation Games
@@
In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo.
@@
(1) Drone Navigation Game. As shown in Fig. 15, the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle) whose position is unknown, reaching different targets (shown in yellow points behind the gate). The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$. The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to specific correction commands to the drone’s control space. See the detailed task setup and keyboard interface in Appendix B-C.
@@
(2) Franka arm reaching game. As in Fig. 16, the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an unknown narrow slot (two yellow bars). Note that the safety constraint is only applied on the end effector and the collision between the link and the slot is ignored. The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis. Human corrections are applied via keyboard, with specific keys mapping to specific correction commands applied to the robot control space $\mathbb{R}^{6}$ (we use operational space control). See the detailed task setup and keyboard interface in Appendix B-D.
@@
In both games, the robots are controlled with MPC controller (3) but fail to maintain safety with the initial safety constraints. To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.
(a) Tracking View
(b) Fixed View
Figure 15: Drone navigation game. (a) Tracking view. (b) Fixed view. Users can see the motion of the drone with the help of these two views. The $xyz$-axis (red, green, and blue axes), the gate (yellow bars), the drone, the local plan solved from the current MPC (red curve) and the motion target (yellow balls) are visible to users. In each game reset, the user will see one target randomly.
(a) Starting Configuration
(b) Motion Illustration
Figure 16: Franka arm reaching game. (a) The starting pose of the arm. RGB arrows denote the world $xyz$-frame. (b) One example of successful reaching, and long green arrow denotes the motion direction. Users can see the motion of the robot arm by adjusting the view. The slot, the robot arm, and current motion target (cyan ball) are all visible to users.
@@ VII-B1 Participants
We recruited 15 participants from Arizona State University, including undergraduates, masters, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are new to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
@@ VII-B2 Procedure
Each participant was instructed to play the previously developed two simulation games. In each game, each participant plays 10 trials.
In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up.
In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets with the end-effector having no collision with the slot. The failure of one trial is determined when there are too many corrections (exceeding 25) or the user believes the robot cannot be corrected anymore and gives up.
The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances to interact with the environment hands-on to get familiar with the keys.
@@ VII-B3 Metrics for Evaluation
Two major metrics are used to evaluate the performance of the proposed algorithm:
  • (I)
    Success rate: The overall and individual success rate suggest the effectiveness of the proposed method.
  • (II)
    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by users in successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better human data efficiency.
@@ VII-C1 User Study Metrics Report
For the user study, we report the success rate and correction count in a successful game trial in Table II. Both metrics are averaged across users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching game. In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses. This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
The above result illustrates that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials). It also shows the efficiency of the proposed method — successful learning only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs. In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be because the drone game has a larger motion range and longer task rollout horizon. As a result, human correction strategies, such as correction position and timing, can vary significantly. This correction diversity leads to a larger variance of correction count.
VII-C2 Analysis for Learned Constraint and Robot Motion
To illustrate learning progress, in Fig. 17 and Fig. 18, we show the intermediate robot motion at different numbers of human corrections. Fig. 17 shows the trajectory of the drone after different numbers of human corrections. Fig. 18 shows the pose of the robot gripper when passing the slot after a different number of human corrections.
@@ VII-C2 Analysis for Learned Constraint and Robot Motion
Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball). (a): After 4 corrections, the gripper collides with the slot. (b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
To better understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned. Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$ (true gate position is $x_{\text{gate}}=9.8$). In the heat map, we draw 00-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary. Also, we plot the $yz$-plane projection of the gate, which is the black box.
Fig. 19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained by the following reasons. As shown in Fig. 15, the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom edge untouched. It is expected that to learn a safety constraint that fully represents the gate, lower target positions should be added so that the drone will receive human corrections near the bottom edge of the gate.
Figure 19: Heatmap for the learned constraint in the drone game. The black solid line denotes the actual position of the gate in $yz$-plane. The black dashed curve denotes the boundary of the learned safety zone. Since the targets are higher than the gate (see Fig. 15), during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
VIII User Study in Real-world Experiment
In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm in a hardware Franka Research 3 experiment. We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to targets. In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns. The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling. The Franka robot arm will learn a safe MPC controller from human physical corrections.
@@ VIII-A Hardware Setup
The setup of mobile water pouring task is shown in Fig. 20. A water pot is gripped by the robot end-effector. The robot moves from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is placed in the middle of the robot path, but the precise size and location of the bar sink are unknown.
Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes connected via arrows that indicate communication via ROS.
As shown Fig. 21, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and safe MPC alignment module. The safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs). The parameters of this Safe MPC controller are listed in Table III. The safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generate the action $\boldsymbol{u}$ (see (58) in Appendix), which is the desired velocity of the end-effector. The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
The safe MPC alignment module updates the parameterize $\boldsymbol{\theta}$ of the safe MPC controller. The update only happens when the alignment module detects a human correction $\boldsymbol{a}$. The module implements our safe MPC alignment algorithm. Note that the alignment module and controller modules runs different process, thus the learning update does not affect the real-time MPC control of the robot arm. ROS is used for communication between different processes.
@@
In the safe MPC controller, the cost of the MPC controller is set to penalize the current and target, as in Section B-D. Since it is difficult to explicitly define a parameterized safety constraint using the state of pouring water, we instead define the constraints only using the robot state. Specifically, we use the weighted RBFs to parameterize the safety constraints constraint, similar to Appendix B-D.
@@
Figure 22: Illustrations of one trial of the real-world experiment. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
@@ VIII-B Participants, Procedures, and Metrics
All participants recruited for the previous two simulation games in Section VII-B also participate the hardware experiment. They are told of the objective before starting trials. One episode is defined as one robot motion from the starting position to the targets, during which the human can apply corrections at any time instances. One learning trial may include multiple episodes (with a maximum of 15 corrections). During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion. The success learning trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VII-B are recorded.

### v3

@@ VI-D Ablation Study
We conduct two ablation studies to show the performance of our algorithm given different design aspects. The task is still based on the pendulum task in Section VI-A.
@@ VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
We compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$: MVE centering (used in our method) and random sampling. For the latter, $\boldsymbol{\theta}_{i}$ is sampled from the uniform distribution supported by $\Theta_{i-1}$. We compare the total number of corrections at learning convergence under the two strategies.
With 50 independent runs, we observe the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling $11.88\pm 3.50$ corrections. While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance. This is understandable given sampling randomness. We also observed that the sampling process introduced a large computational overhead, which is around 25% time slower than MVE centering. In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency. Because selecting $\boldsymbol{\theta}_{i}$ is cast as a convex programming for MVE centering, such an advantage would be more significant for a high-dimensional parameter space.
@@ VI-D2 Choosing $\gamma$
[…]
Figure 15: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale.
@@ VII User Study in Simulation Games
We conduct two human user studies to evaluate the Safe MPC Alignment. One is the drone navigation game and the other the Franka robot arm reaching game. In both games, the robot is controlled by an MPC policy and human participants provide real-time corrections through a keyboard. The objective of each game is to teach a robot to successfully learn Safe MPC to move through environments with obstacles. Section VII-A describes the settings. Section VII-B details the user study procedure. The results are given in Section VII-C.
VII-A Two Simulation Games
@@
In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo [todorov2012mujoco].
@@
(1) Drone Navigation Game. As shown in Fig. 16(a), the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle), reaching different targets (shown in yellow points behind the gate). The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$. The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to the drone’s control space. See the details in Appendix B-C.
@@
(2) Franka arm reaching game. As shown in Fig. 16(b), the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an narrow slot (two yellow bars). Note that the safety constraint is only applied to the end effector and the collision between the link and the slot is ignored. The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis. Human corrections are applied via keyboard, with specific keys mapping to the robot control space $\mathbb{R}^{6}$ (we use operational space control). See the details in Appendix B-D.
@@
In both games, the robots are controlled with Safe MPC policy (3) but fail to maintain safety with the initial safety constraints. To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.
(a) Drone Navigation game
(b) Franka arm reaching game
Figure 16: Two simulated games in user study. (a) Drone Navigation game. (b) Franka arm reaching game.
@@ VII-B1 Participants
We recruited 15 participants from Arizona State University, including undergraduates, masters, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are new to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB), and all participants had signed the consent forms.
@@ VII-B2 Procedure
Each participant was asked to play the two simulation games. In each game, each participant plays 10 trials. In the drone navigation game, a trial is successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up. In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets with the end-effector having no collision with the slot. The failure of one trial is determined when there are too many corrections (exceeding 25) or the user believes the robot cannot be corrected anymore and gives up. The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances to interact with the environment hands-on to get familiar with the keys.
@@ VII-B3 Metrics for Evaluation
Two metrics are used to evaluate the outcomes: (I) Success rate: The overall success rate suggest the effectiveness of the proposed method. (II) The number of corrections needed for a successful trial: The averaged number of user corrections for successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better human data efficiency.
@@ VII-C1 Results
We report the success rate and correction count in a successful game trial in Table II. Both metrics are averaged across users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ for Franka arm reaching game. In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses. This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
The above result shows that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials). It also shows the efficiency of the proposed method: successful learning only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs. In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be because the drone game has a larger motion range and longer rollout horizon. As a result, human correction strategies, such as correction position and timing, vary more.
VII-C2 Analysis
To illustrate learning progress, in Fig. 17 and Fig. 18, we show the intermediate robot motion after different numbers of human corrections. Fig. 17 shows the safe MPC rollout trajectory of the drone after different numbers of human corrections. Fig. 18 shows the keyframes of the robot gripper when passing the slot after a different human corrections.
@@ VII-C2 Analysis
Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball in different columns). (a): After 4 corrections, the gripper collides with the slot. (b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
To understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned. Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$. In the heat map, we draw 0-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary. Also, we plot y-z plane projection of the gate, which is the black box.
Fig. 19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained below. As shown in Fig. 16(a), the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 17. As a result, the majority of human corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom edge untouched. It is expected that to learn a safety constraint that fully represents the gate, lower target positions should be added so that the drone will receive human corrections near the bottom edge of the gate.
Figure 19: Heatmap for the learned constraint in the drone game. The black solid line denotes the actual position of the gate in y-z plane. The black dashed curve denotes the boundary of the learned safety zone. Since the majority of human directional corrections happen when the drone approaches the top edge of the gate, the learned safety boundary is accurate near the top edge of the gate and the bottom boundary is untouched.
VIII Real-world Experiment
We conduct a user study to evaluate the performance of the Safe MPC Alignment in a real-world experiment. We focus on the task of mobile liquid pouring: the Franka robot arm performs liquid pouring while moving from one location to targets. The task is challenging because it is hard to manually define a safety constraint to accomplish liquid pouring without spilling. The robot arm will learn a safe MPC controller from human physical corrections.
@@ VIII-A Hardware Setup
The setup of the mobile water pouring task is shown in Fig. 20. A water pot is gripped by the robot’s end-effector. The robot moves from the starting point $0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is placed in the middle of the robot’s path, but the precise size and location of the bar sink are unknown.
Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes communicated via ROS.
Fig. 21 shows three main components of the system: the parameterized Safe MPC controller (high-level controller), the low-level controller, and the Safe MPC alignment module. The Safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs). The parameters of this Safe MPC controller are listed in Table III. The Safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generates the action $\boldsymbol{u}$ (see (59) in Appendix), which is the desired velocity of the end-effector. The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
The Safe MPC Alignment module updates the parameter $\boldsymbol{\theta}$ of the Safe MPC controller. The update only happens when the alignment module detects a human correction $\boldsymbol{a}$. The module implements our Safe MPC Alignment algorithm. Note that the alignment module and controller modules runs in different processes, thus the learning update does not affect the real-time MPC control of the robot arm. ROS is used for communication between different processes.
@@
Figure 22: Illustrations of one trial of the real-world experiment. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and the altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x$-axis and successfully aligns with human intent.
@@ VIII-B Participants, Procedures, and Metrics
All participants recruited in Section VII-B also participate in the hardware experiment. They are told of the objective before starting trials. One episode is defined as one robot motion from the starting position to the targets, during which the human can apply corrections at any time instances. One learning trial may include multiple episodes (with a maximum of 15 corrections). During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot. The successful trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VII-B are recorded.